OpenStudy (anonymous):

@iGreen What is the approximate volume of a cylinder with a diameter of 10 meters and a height of 5 meters? Use 3.14 for pi. A. 130.8 m^3 B. 314 m^3 C. 1570 m^3 D. 392.5 m^3

OpenStudy (january123):

= (3.14)(20/2)²(5) = 15.7(10)² = 15.7(100) = 1,570

OpenStudy (anonymous):

v=3.14xr2xh

OpenStudy (anonymous):

@january123 YOU DON'T HAVE TO GIVE THE ANSWERS

OpenStudy (january123):

Well, all i did was google it and c\p it on here and so yeah don't get mad dude

OpenStudy (shamim):

Can u tell me Cross sectional area of ur given cylinder A=?

OpenStudy (anonymous):

I'm not mad, just sayin' ^_^

OpenStudy (shamim):

U know A=pi*r^2

OpenStudy (shamim):

U r voilating rules of openstudy @january123

OpenStudy (january123):

eh what EVER, CONSIDER ME OUT OF THIS QUESTION BECAUSE PPL R JUST CRYING OVER STUPID THINGS, SO BIE PIE

OpenStudy (january123):

@Master_Chief_117 post a reply so i can give u a medal